# On the Vulnerability of Applying Retrieval-Augmented Generation within Knowledge-Intensive Application Domains

### v1

@@ 4 Understanding the vulnerability of retrieval systems in RAG
In this section, we provide insights towards understanding the vulnerability for the retrieval systems. To begin with, we will first present an intriguing property shared by popular retrievers, which we termed as the orthogonal augmentation property. The orthogonal augmentation property states that when two documents $q,p\in\mathcal{V}^{L}$ are close to orthogonal in terms of their embedding, the embedding of the (token-level) concatenated one, i.e., $f([q\oplus p])$, roughly equals $f(q)+v$, with $v^{\mathrm{\scriptscriptstyle T}}f(q)\approx 0$. In other words, this property implies that the shift (in terms of embeddings) from $q$ to $[q\oplus p]$ mainly occurs in directions that are perpendicular to $q$. As a result, for the inner-product based similarity, the similarity between $q$ and $[q\oplus p]$ will roughly equal to that between $q$ and itself, namely $f(q)^{\mathrm{\scriptscriptstyle T}}f([q\oplus p])\approx f(q)^{\mathrm{\scriptscriptstyle T}}f(q)+f(q)^{\mathrm{\scriptscriptstyle T}}v\approx f(q)^{\mathrm{\scriptscriptstyle T}}f(q)$. This implies that $[q\oplus p]$ is likely to be retrieved by $q$, possibly with high ranking, indicating the success of universal poisoning attacks.
@@
We present the similarity measurements between the query and the clean retrieved documents (from the corpus Textbook) in Table 4 for Contriever and MedCPT, respectively. For each query, we calculate its similarity, both the cosine similarity and the inner product, with the $K=5$-nearest neighbor retrieved documents from the corpus Textbook, and report the mean and standard deviation. Additionally, we also report the value of the inner product between a query and itself for the purpose of comparison. We observed that both the inner product and the cosine similarity are low across all query sets. For instance, an average angle of around $70^{\circ}$ between the query and retrieved documents is observed across all query sets for the Contriever, and the inner product between the query and retrieved documents is less than $25\%$ of that between the query and itself. These results indicate that the retrieved documents are not as close to their queries as one might expect. In conclusion, these findings highlight a gap between the query and its clean retrieved documents. This disparity leaves room for exploitation by various adversarial attacks, including our proposed universal poisoning attacks, posing considerable safety risks.
@@
Table 4: Evidence on the low similarity of the (clean) retrieved documents. For each query $q$ sampled from the query set, we calculate its similarity, both in angle (denoted by (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ )) and inner product, with the $K=5$-nearest neighbor retrieved documents from the corpus Textbook, denoted as $R$, and then report the mean and standard deviation. We observed that both the inner product and the cosine similarity are low across all query sets, indicating the low quality of the retrieved documents.
@@ 5 New Defense
In this section, we propose a detection-based method to defend against the proposed universal poisoning attacks. We consider a scenario in which the defender, such as an RAG service provider, has full access to retrievers. They collect documents from public websites, integrate them into their data corpus, and provide services using both the retrievers and the updated data corpus. The defender’s objective is to develop an algorithm capable of automatically detecting potential adversarial documents to be incorporated into their data corpus. Without loss of generality, we assume that the defender already has a collection of clean documents associated with a set of targeted queries to be protected, which serve as an anchor set (denoted as $\mathcal{A}=\{a_{1},\ldots,a_{|\mathcal{A}|}\}$) for detection.
@@
Due to the high-dimensional nature of the embeddings, calculating the Mahalanobis distance can be challenging. This is because the sample covariance matrix ensued can be numerically unstable in large data dimensions [40, 41], leading to an ill-conditioned matrix that is difficult to invert [42]. To address this issue, we consider regularizing the sample covariance matrix through shrinkage techniques [43, 44]. In detail, we conduct shrinkage by shifting each eigenvalue by a certain amount, which in practice leads to the following, $s(X)\triangleq(X-\mu)^{\top}\Sigma_{\beta}^{-1}(X-\mu),$ where $\mu$ is the mean of $\{f(a_{i})\}_{i=1}^{|\mathcal{A}|}$, with $a_{i}\in\mathcal{A}$, ($A$ is the anchor set defined earlier in this section) and $\Sigma_{\beta}\triangleq(1-\beta)S+d^{-1}\beta{\rm Tr}(S)I_{d},$ with $S$ being the sample covariance of $\{f(a_{i})\}_{i=1}^{|\mathcal{A}|}$, ${\rm Tr}(\cdot)$ is the trace operator, and $\beta\in(0,1)$ is the shrinkage level. We select $\beta$ by cross-validation, with an ablation study in Appendix C.2.
We tested the proposed method for filtering out the poisoned documents over the previous setups. The threshold for filtering is set at the $95$th quantile of the $\{s(f(a_{i}))\}_{i=1}^{|\mathcal{A}|}.$ We present some case studies in Figure 3 below. The complete results are included in Figure 4 and Figure 5 for PubMed (in the appendix), Figure 6 and Figure 7 for StatPearl (in the appendix), and Figure 9 and Figure 8 for Textbook (in the appendix). We observed that, in almost all cases, the proposed method achieves near-perfect detection. Additionally, we observed that the $\ell_{2}$-norm defense is effective when using the Contriever. One potential reason may be that it is sensitive to the total length of the documents. Furthermore, we applied our proposed method to one of the state-of-the-art poisoning attacks [15] and obtained a filtering rate greater than 95%, indicating the wide applicability of our proposed defense.
@@ 6 Conclusion
This paper studies the vulnerability of retrieval systems in RAG. We first demonstrate that retrieval systems in RAG for medical Q&A are vulnerable to universal poisoning attacks. Next, we provide two-fold insights towards understanding the vulnerability: (1) by identifying an intriguing property of dense retrievers, and (2) revealing the relatively low similarity of the clean retrieved documents. Based on these findings, we develop a detection-based defense, achieving high detection accuracy.
Limitations & Future Work There are several potential directions that can be further explored. First, the experimental studies are only conducted on medical Q&A. Investigating the use of RAG in other knowledge-intensive application domains to determine if they suffer from similar security risks is important. Second, the empirical results showed that the retrieved clean documents are not close to their associated queries, leaving a large gap that can be exploited for adversarial attacks. An interesting question is whether we can develop methods to align the queries and the documents in the corpus to enhance retrieval quality. Third, although the developed defense can effectively detect a large portion of poisoned documents, its usage is limited to a set of targeted queries instead of arbitrary queries. Another important direction is to extend the current method to arbitrary-query sets.
Boarder Impacts There are both potential positive and negative societal impacts of this work. Potential negative impacts include the possibility that an adversary could apply the proposed attack to other retrieval systems. On the other hand, we anticipate that there will be many more positive societal impacts of this work: (1) highlighting the need for vigilance regarding security risks when applying RAG in safety-critical domains; (2) providing several insights into understanding these potential safety risks; and (3) introducing a new, effective defense that can be used for detecting poisoned documents.
@@ Acknowledgement
In section A, we list more implementation details including: computing resources, examples of datasets and others. We provide experimental results are omitted from the main text due to space limit in Section B. In section C, we provide ablation studies regarding different choices of hyperparameters used in the paper.
@@ A.2.2 Query
Following [14], we provide the statistics of used query in Table 6 below. We also provide some examples of these quetions in the following.

### v2

@@ 4 Understanding the vulnerability of retrieval systems in RAG
In this section, we provide insights towards understanding the vulnerability for the retrieval systems. To begin with, we will first present an intriguing property shared by popular retrievers, which we termed as the orthogonal augmentation property. The orthogonal augmentation property states that when two documents $q,p\in\mathcal{V}^{L}$ are close to orthogonal in terms of their embedding, the embedding of the (token-level) concatenated one, i.e., $f([q\oplus p])$, roughly equals $f(q)+v$, with $v^{\mathrm{\scriptscriptstyle T}}f(q)\approx 0$. In other words, this property implies that the shift (in terms of embeddings) from $q$ to $[q\oplus p]$ mainly occurs in directions that are perpendicular to $q$. As a result, for the inner-product based similarity, the similarity between $q$ and $[q\oplus p]$ will roughly equal to that between $q$ and itself, namely $f(q)^{\mathrm{\scriptscriptstyle T}}f([q\oplus p])\approx f(q)^{\mathrm{\scriptscriptstyle T}}f(q)+f(q)^{\mathrm{\scriptscriptstyle T}}v\approx f(q)^{\mathrm{\scriptscriptstyle T}}f(q)$. This implies that $[q\oplus p]$ is likely to be retrieved by $q$, indicating the success of universal poisoning attacks.
@@
We present similarity measurements between the query and the clean retrieved documents (from the corpus Textbook) in Table 4 for Contriever and MedCPT, respectively. For each query, we calculate its similarity, the cosine similarity and the inner product, with the $K=5$-nearest neighbor retrieved documents from the corpus Textbook, and report the mean and standard deviation. Additionally, we also report the value of the inner product between a query and itself for comparison. We observed that both the inner product and the cosine similarity are low across all query sets. For instance, an average angle of around $70^{\circ}$ between the query and retrieved documents is observed across all query sets for the Contriever, and the inner product between the query and retrieved documents is less than $25\%$ of that between the query and itself. These results indicate that the retrieved documents are not as close to their queries as one might expect. In conclusion, these findings highlight a gap between the query and its clean retrieved documents. This disparity leaves room for exploitation by various adversarial attacks, including our proposed universal poisoning attacks, posing considerable safety risks.
@@
Table 4: Evidence on the low similarity of the (clean) retrieved documents. For each query $q$ sampled from the query set, we calculate its similarity, both in angle (denoted by (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ )) and inner product, with the $5$-nearest neighbor retrieved documents from the corpus Textbook, denoted as $R$, and then report the mean and standard deviation. We observed that both the inner product and the cosine similarity are low across all query sets, indicating the low quality of the retrieved documents.
@@ 5 New Defense
In this section, we propose a detection-based method to defend against the proposed universal poisoning attacks. We consider a scenario in which the defender, such as an RAG service provider, has full access to retrievers. They collect documents from public websites, integrate them into their data corpus, and provide services using both the retrievers and the updated data corpus. The defender’s objective is to develop an algorithm capable of automatically detecting potential adversarial documents to be incorporated into their data corpus. We assume that the defender already has a collection of clean documents associated with a set of targeted queries to be protected, which serve as an anchor set (denoted as $\mathcal{A}=\{a_{1},\ldots,a_{|\mathcal{A}|}\}$) for detection.
@@
Due to the high dimension of embeddings, calculating the Mahalanobis distance can be challenging. The sample covariance matrix ensued can be numerically unstable in large data dimensions [46], leading to an ill-conditioned matrix that is difficult to invert [47]. To address this issue, we consider regularizing the sample covariance matrix through shrinkage techniques [48, 49]. In detail, we conduct shrinkage by shifting each eigenvalue by a certain amount, which in practice leads to the following, $s(X)\triangleq(X-\mu)^{\top}\Sigma_{\beta}^{-1}(X-\mu),$ where $\mu$ is the mean of $\{f(a_{i})\}_{i=1}^{|\mathcal{A}|}$, with $a_{i}\in\mathcal{A}$ (the anchor set defined earlier), and $\Sigma_{\beta}\triangleq(1-\beta)S+d^{-1}\beta{\rm Tr}(S)I_{d},$ with $S$ being the sample covariance of $\{f(a_{i})\}_{i=1}^{|\mathcal{A}|}$, ${\rm Tr}(\cdot)$ is the trace operator, and $\beta\in(0,1)$ is the shrinkage level. We select $\beta$ by cross-validation, with an ablation study in Appendix C.2.
We tested the proposed method for filtering out the poisoned documents over the previous setups. The threshold for filtering is set at the $95$th quantile of the $\{s(f(a_{i}))\}_{i=1}^{|\mathcal{A}|}.$ We present some case studies in Figure 3 below. The complete results are included in Figure 4 and 5 for PubMed (in the appendix), Figure 6 and 7 for StatPearl (in the appendix), and Figure 8 and 9 for Textbook (in the appendix). We observed that, in almost all cases, the proposed method achieves near-perfect detection. Additionally, we observed that the $\ell_{2}$-norm defense is effective when using the Contriever. One potential reason may be that it is sensitive to the total length of the documents. Furthermore, we applied our method to one of the state-of-the-art poisoning attacks [15] and obtained a filtering rate greater than 95%, indicating the wide applicability of our proposed defense.
@@ 6 Conclusion
This paper studies the vulnerability of retrieval systems in RAG. We first demonstrate that retrieval systems in RAG for medical Q&A are vulnerable to universal poisoning attacks. Next, we provide two-fold insights towards understanding the vulnerability: (1) by identifying an intriguing property of dense retrievers, and (2) revealing the relatively low similarity of the clean retrieved documents. Based on these findings, we develop a detection-based defense, achieving high detection accuracy. Recent work on dynamic pruning for LLMs, such as Probe Pruning [50], enables efficient inference without model retraining. Investigating how such pruning methods affect the robustness and retrieval behavior of RAG systems remains an important direction for future work. The Appendix provides additional details on experimental setups and ablation studies.
Impact Statement
This paper presents work aimed at advancing the safety and trustworthiness of machine learning. There are both potential positive and negative societal impacts of this work. Potential negative impacts include the possibility that an adversary could apply the proposed attack to other retrieval systems. On the other hand, we anticipate that there will be many more positive societal impacts of this work: (1) highlighting the need for vigilance regarding security risks when applying RAG in safety-critical domains; (2) providing several insights into understanding these potential safety risks; and (3) introducing a new, effective defense that can be used for detecting poisoned documents.
@@ Acknowledgement
In section A, we list more implementation details including: computing resources, examples of datasets and others. We provide experimental results are omitted from the main text due to space limit in Section B. In section C, we provide ablation studies regarding different choices of hyperparameters used in the paper. In Section D, we present an empirical study on applying our proposed attacks to legal domains.
@@ A.2.2 Query
Following [14], we provide the statistics of used query in Table 6 below. We also provide some examples of these questions in the following.
@@ Appendix B Omitted experimental results
Performance under paraphrasing defenses The table below presents the results for the proposed attacks under paraphrasing defenses, where the defender rephrases the queries as a defense mechanism to prevent the attack from knowing the exact queries. We observe only a slight decrease in attack success rates compared to the case using exact queries, highlighting the robustness of the proposed attacks.
Corpus Retriever Query Target Information
PPI MS-MARCO NQ HotpotQA Diagnostic
Textbook MedCPT MMLU $0.76$ $0.72$ $0.61$ $0.52$ $0.64$
MedQA $0.98$ $0.97$ $0.98$ $0.97$ $0.96$
PubMedQA $0.94$ $0.96$ $0.92$ $0.95$ $0.91$
Specter MMLU $0.79$ $0.75$ $0.54$ $0.84$ $0.72$
MedQA $0.97$ $0.86$ $0.87$ $0.97$ $0.91$
PubMedQA $0.94$ $0.80$ $0.85$ $0.98$ $0.82$
Contriever MMLU $0.57$ $0.75$ $0.64$ $0.85$ $0.70$
MedQA $0.90$ $0.91$ $0.94$ $0.96$ $0.90$
PubMedQA $0.97$ $0.98$ $0.98$ $0.98$ $0.92$
@@ C.2 On the shrinkage level $\beta$
Appendix D Examples of Legal Q&A Datasets
In this section, we list some examples of the legal Q&A dataset used in our paper in the table below.
Table 10: Examples Legal Q&A case
Component Description
Case Name Smith v The State of New South Wales (2015)
Question How did the case of Smith v The State of New South Wales
(2015) clarify the interpretation of ‘reasonable grounds’ for
conducting a search without a warrant?
Support In [Case Name: Smith v The State of NSW (2015)],
the plaintiff [Action: was searched without a warrant]
[Location: near a known drug trafficking area] based
on the plaintiff’s nervous demeanor and presence in the
area, but [Outcome: no drugs were found]. The legality
of the search was contested, focusing on [Legal Concept:
whether ‘reasonable grounds’ existed].
Answer The case ruled ‘reasonable grounds’ require clear, specific
facts of criminal activity, not just location or behavior.